OpenStudy (anonymous):

-6 l2x-14l=-42 absolute value.

OpenStudy (saifoo.khan):

|2x - 14| = 7

OpenStudy (saifoo.khan):

2x - 14 = 7 & 2x - 14 =-7

OpenStudy (anonymous):

|2x - 14| = 7 2x - 14 = 7 x = 21/2 x = 10.5 or 2x - 14 = -7 x = 7/2 x = 3.5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

-6 l2x-14l=-42 ===> l2x-14l=7 ===> 2x-14=7 ==>x=21/2 , or 2x-14=-7 ==> x=7/2